# Exact Random Graph Matching with Multiple Graphs

### v1

@@ Appendix A Maximum Likelihood Estimator
i.e. maximizing the likelihood corresponds to selecting $\pi_{12},\cdots,\pi_{1m}$ to maximize $\alpha_{0}$ the number of node pairs $e$ for which $G_{1}(e)=G_{2}(\pi_{12}(e))=\cdots=G_{m}(\pi_{1m}(e))=0$. This is equivalent to minimizing the number of edges in the union graph $G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}}$, as desired. ∎
Remark 12.
In the case of two graphs, minimizing the number of edges in the union graph $G_{1}\vee_{\pi}G_{2}$ is equivalent to maximizing the number of edges in the intersection graph $G_{1}\wedge_{\pi}G_{2}$. This is consistent with existing literature on two graphs [CK16, WXY22].
Appendix B Concentration Inequalities for Binomial Random Variables
The following bounds for the binomial distribution are used frequently in the analysis.
Lemma 13.
Let $X\sim\mathsf{Bin}(n,p)$. Then,
  1. 1.
    For any $\delta>0$,
  2. 2.
    For any $\delta>5$,
  3. 3.
    For any $\delta\in(0,1)$,
    $\mathbb{P}\left(X\leq(1-\delta)np\right)\leq\left(\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right)^{np}.$ (7)
Proof.
All proofs follow from the Chernoff bound and can be found, or easily derived, from Theorems 4.4 and 4.5 of [MU17]. ∎
Appendix C Proof of Lemma 8
We restate Lemma 8 for convenience. See 8
Before presenting the proof, we present the intuition behind it. The events $\{v\notin\mathsf{core}_{k}(G)\}$ and $\left\{\delta_{G}(v)\geq k+1/\alpha\right\}$ are highly negatively correlated. However, consider the subgraph $(G-v)$ of $G$ induced on the vertex set $V-\{v\}$, and note that the $k$-core of this subgraph does not depend on the degree of $v$. Furthermore, if $v\notin\mathsf{core}_{k}(G)$, then it must be that $v$ has fewer than $k$ neighbors in $\mathsf{core}_{k}(G-v)$. Intuitively, this event has low probability if $\mathsf{core}_{k}(G-v)$ is sufficiently large.
Notice that $(G-v)\sim\mathsf{ER}(n-1,\alpha\log(n)/n)$, and so standard results about the size of the $k$-core of Erdős-Rényi graphs apply. However, we require the error probability that the $k$-core of $G-v$ is too small to be $o(1/n)$ this is crucial since we will later use a union bound over all the nodes $v$. Unfortunately, standard results such as [Łuc91] can only be invoked directly to show that the corresponding probability is $o(1)$, which is insufficient for our purpose. Later in this section, we refine the analysis in [Łuc91] to obtain the desired convergence rate. The refinement culminates in the following.
Lemma 14.
Let $\alpha>0$ and $G\sim\mathsf{ER}(n,\alpha\log(n)/n)$. Let $v$ be a node of $G$. The size of the $k$-core of $G-v$ satisfies
$\mathbb{P}\left(\left|\mathsf{core}_{k}(G-v)\right|<n-3n^{1-\alpha}\right)=o(1/n).$
The proof of Lemma 14 is deferred to Section C.1. It remains to study the error event that $v$ has too few neighbors in $\mathsf{core}_{k}(G-v)$. To count the number of neighbors of $v$ in $\mathsf{core}_{k}(G-v)$, we exploit the independence of $\mathsf{core}_{k}(G-v)$ and $v$ as follows: each neighbor of $v$ is considered a success if it belongs to $\mathsf{core}_{k}(G-v)$ and a failure otherwise. Counting the number of successes is equivalent to sampling with replacement $\delta_{G}(v)$ elements, each of which is independently a success with probability $|\mathsf{core}_{k}(G-v)|/\left(n-1\right)$. The number of successes then follows precisely a hypergeometric distribution. This intuition is made rigorous in the proof below.
Let us recall some facts about the hypergeometric distribution because it plays an important role in the proof. Denote by $\mathsf{HypGeom}(N,K,n)$ a random variable that counts the number of successes in a sample of $n$ elements drawn without replacement from a population of $N$ individuals, of which $K$ elements are considered successes. Note that if this sampling were done with replacement, then the number of successes would follow a $\mathsf{Bin}\left(n,K/N\right)$ distribution. A result of Hoeffding [Hoe94] establishes that the $\mathsf{HypGeom}(N,K,n)$ distribution is convex-order dominated by the $\mathsf{Bin}\left(n,K/N\right)$ distribution, i.e.
$\mathbb{E}\left[f(\mathsf{HypGeom}(N,K,n))\right]\leq\mathbb{E}\left[f(\mathsf{Bin}(n,K/N))\right]\ \text{ for all convex functions $f$}.$
In particular, the function $f(x)=e^{tx}$ is convex for any value of $t$, and so Chernoff bounds that hold for the binomial distribution also hold for the corresponding hypergeometric distribution. This yields the following proposition.
Proposition 15.
Let $X\sim\mathsf{HypGeom}(N,K,n)$. It follows for any $\delta>0$ that
$\mathbb{P}\left(X>(1+\delta)\times\frac{nK}{N}\right)\leq\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{nK/N}\leq\left(\frac{e}{1+\delta}\right)^{(1+\delta)nK/N}.$
Our final remark about the hypergeometric distribution is a symmetry property. By interchanging the success and failure states, it follows that
$\mathbb{P}\left(\mathsf{HypGeom}(N,K,n)=k\right)=\mathbb{P}\left(\mathsf{HypGeom}(N,N-K,n)=n-k\right).$
The above intuition for the proof of Lemma 8 is formalized below.
Proof of Lemma 8. Let $V$ denote the vertex set of $G$, and let $G-v$ denote the induced subgraph of $G$ on the vertex set $V-\{v\}$. For any set $A\subseteq V$, let $N_{v}(A)$ denote the set of neighbors of $v$ in the set $A$, i.e.
$N_{v}(A):=\left\{u\in A:\{u,v\}\in E(G)\right\}.$
Since $\mathsf{core}_{k}(G-v)\subseteq\mathsf{core}_{k}(G)$, it is true that
$\left\{v\notin\mathsf{core}_{k}(G)\right\}\subseteq\left\{|N_{v}(\mathsf{core}_{k}(G))|\leq k-1\right\}\subseteq\left\{|N_{v}(\mathsf{core}_{k}(G-v))|\leq k-1\right\}.$
It follows that
($\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\right)\leq p_{1}+p_{2},$
where
$p_{1}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|<n-3n^{1-\alpha}\right\}\right)$
$p_{2}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right\}\right)$
It suffices to show that both $p_{1}$ and $p_{2}$ are $o(1/n)$. The term $p_{1}$ deals with the probability that the $k$-core of $G-v$ is too small. In fact, by Lemma 14, it follows directly that
$p_{1}$ $\leq\mathbb{P}\left(|\mathsf{core}_{k}(G-v)|<n-3n^{1-\alpha}\right)=o(1/n),$
Next, the probability $p_{2}$ is analyzed. Enumerate arbitrarily but independently the elements of sets $N_{v}(V)$ and $\mathsf{core}_{k}(G-v)$, so that
$N_{v}(V)=\left\{v_{1},\cdots,v_{\delta_{G}(v)}\right\},\quad\mathsf{core}_{k}(G-v)=\left\{a_{1},\cdots,a_{|\mathsf{core}_{k}(G-v)|}\right\}.$
Given that $N_{v}(V)$ has more than $k+1/\alpha$ nodes and $\mathsf{core}_{k}(G-v)$ has more than $n-3n^{1-\alpha}$ nodes, it is true that
$N_{v}(\mathsf{core}_{k}(G-v))$ $=N_{v}(V)\cap\mathsf{core}_{k}(G-v)$
$\supseteq\left\{v_{1},\cdots,v_{\lceil k+1/\alpha\rceil}\right\}\cap\left\{a_{1},\cdots,a_{\lceil n-3n^{1-\alpha}\rceil}\right\}=:\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v)).$
In words, $\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))$ counts among the first $\lceil k+1/\alpha\rceil$ neighbors of $v$ those nodes that are also in the first $\lceil n-3n^{1-\alpha}\rceil$ nodes of $\mathsf{core}_{k}(G-v)$. Therefore,
$p_{2}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq\ k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right\}\right)$
($\leq\mathbb{P}\left(\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\}\ \big{|}\ \delta_{G}(v)\geq\ k+1/\alpha,\ |\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right)$
($\leq\mathbb{P}\left(\{\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))\leq k-1\}\ \big{|}\ \delta_{G}(v)\geq\ k+1/\alpha,\ |\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right)$ (8)
Note that $\mathsf{core}_{k}(G-v)$ is entirely determined by the graph $G-v$, i.e. it is independent of the neighbors of $v$. Consequently, the two sets $\left\{v_{1},\cdots,v_{\lceil k+1/\alpha\rceil}\right\}$ and $\{a_{1},\cdots,a_{\lceil n-3n^{1-\alpha}\rceil}\}$ are selected independent of each other. Equivalently, given that $|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}$ and $\delta_{G}(v)\geq k+1/\alpha$, the size of the intersection set $\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))$ follows a hypergeometric distribution with parameters $(n-1,\lceil n-3n^{1-\alpha}\rceil,\lceil k+1/\alpha\rceil)$. Therefore,
@@ Appendix C Proof of Lemma 8
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathbb{P}\left(\mathsf{HypGeom}(n-1,n-1-\lceil n-3n^{1-\alpha}\rceil,\lceil k+1/\alpha\rceil)\geq\lceil k+1/\alpha\rceil-(k-1)\right)$
$=\mathbb{P}\left(\mathsf{HypGeom}(n-1,\lfloor 3n^{1-\alpha}\rfloor-1,\lceil k+1/\alpha\rceil)\geq 1+1/\alpha\right)$ (9)
where (a) uses the symmetry of the hypergeometric distribution. Using Proposition 15 and the fact that $n-1\geq n/2$ for any $n\geq 1$ yields
@@
$\leq\left(\frac{6e\lceil k+1/\alpha\rceil}{1+1/\alpha}\right)^{1+1/\alpha}\times n^{-1-\alpha}$
$=o(1/n),$
whenever $\alpha>0$. ∎
C.1 Proof of Lemma 14
A key ingredient towards proving Lemma 14 is a useful result about the number of low-degree vertices in an Erdős-Rényi graph, presented next.
Proposition 16.
Let $\alpha>0$ and $G\sim\mathsf{ER}\left(n-1,\alpha\log(n)/n\right)$. Let $r$ be a positive integer and let $Z_{r}$ denote the set of vertices in $G$ with degree no more than $r$, i.e.
$Z_{r}=\left\{v\in V(G):\delta_{G}(v)\leq r\right\}.$
For any $\delta$ such that $\delta>1-\alpha$, it is true that
$\mathbb{P}\left(|Z_{r}|\geq n^{\delta}\right)=o(1/n).$
Proof.
Notice that
$\mathbb{P}\left(|Z_{r}|\geq n^{\delta}\right)$ $=\mathbb{P}\left(\exists S^{\prime}\subseteq V:\left\{|S^{\prime}|\geq n^{\delta}\right\}\cap\left\{\max_{i\in S^{\prime}}\delta_{G}(i)\leq r\right\}\right)$
$\leq\mathbb{P}\left(\exists\ S\subseteq V:\ \left\{|S|=n^{\delta}\right\}\cap\left\{\max_{i\in S}\delta_{G}(i)\leq r\right\}\right)$
$\leq\mathbb{P}\left(\exists\ S\subseteq V:\ \left\{|S|=n^{\delta}\right\}\cap\left\{\sum_{i\in S}\delta_{G}(i)\leq r|S|\right\}\right).$ (10)
If $|S|=n^{\delta}$, then the sum of degrees of vertices in $S$ is the total number of edges with exactly end point in $S$, plus twice the number of edges with both end points in $S$. There are exactly $\binom{|S|}{2}+|S|\left(n-1-|S|\right)\leq n^{1+\delta}$ such vertex pairs, and each of them independently has an edge with probability $\alpha\log(n)/n$. Therefore, a union bound over all possible choices of $S$ yields

### v2

@@ Appendix A Maximum Likelihood Estimator for Exactly Matching Multiple Homogeneous Graphs
Remark 28
In the case of two graphs, minimizing the number of edges in the union graph $G_{1}\vee_{\pi}G_{2}$ is equivalent to maximizing the number of edges in the intersection graph $G_{1}\wedge_{\pi}G_{2}$. This is consistent with existing literature on two graphs [5, 45].
Appendix B On negative results for exact recovery with two graphs
We build up to the proof of Lemma 9 by presenting first a useful lemma about inhomogeneous random graphs.
Lemma 29
Let $G\sim\mathrm{RG}(n,\mathbf{Q})$ be an inhomogeneous random graph, and let $N\triangleq\binom{K}{2}$, where $K$ is the number of isolated nodes in $G$. Then,
$\mathbb{P}\left(N\geq 1\right)\geq\exp(-(2n-4)p_{\max}^{2}-6p_{\max})\cdot\frac{\mu^{2}-2\mu\max_{1\leq i\leq n}e^{-d_{i}}}{4(\mu^{2}+\mu+1)},$ (32)
where $\mu\triangleq\sum_{i=1}^{n}e^{-d_{i}}$.
Proof. Let $Z_{i}$ denote the indicator random variable that node $i$ is isolated in $G$. Further, let $\widetilde{Z}_{i}\sim\mathrm{Bern}(e^{-d_{i}})$ denote a collection of independent random variables and define $\widetilde{N}\triangleq\sum_{1\leq i<j\leq n}\widetilde{Z}_{i}\cdot\widetilde{Z}_{j}$. We use the second moment method by relating the moments of $N$ and $\widetilde{N}$.
Bounding the second moment. Notice that
$\mathbb{E}[N^{2}]$ $=\sum_{1\leq i<j\leq n}\ \sum_{1\leq k<\ell\leq n}\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]\,,$ (33)
$\mathbb{E}[\widetilde{N}^{2}]$ $=\sum_{1\leq i<j\leq n}\ \sum_{1\leq k<\ell\leq n}\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}].$ (34)
Consider the following three exhaustive cases.
  • Case $1$: $i=k$ and $j=\ell$. Notice that
    $\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]=\mathbb{E}[\widetilde{Z}_{i}^{2}]\cdot\mathbb{E}[\widetilde{Z}_{j}^{2}]=e^{-d_{i}-d_{j}}.$
    Therefore,
    $\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]\!=\!(1-p_{ij})\prod_{\begin{subarray}{c}{(i,u)}\\ u\neq j\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}{(j,v)}\\ v\neq i\end{subarray}}(1-p_{jv})\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}e^{-d_{i}-d_{j}+p_{ij}}\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}],$ (35)
    where (a) uses that $1-x\leq e^{-x}$ for all $x$, and (b) uses that $p_{ij}\leq p_{\max}\leq 6p_{\max}$.
  • Case $2$: $j=k$. Notice that
    $\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]=\mathbb{E}[\widetilde{Z}_{i}]\cdot\mathbb{E}[\widetilde{Z}_{j}^{2}]\cdot\mathbb{E}[\widetilde{Z}_{\ell}]=e^{-d_{i}-d_{j}-d_{\ell}}.$
    Therefore, similar to case $1$, we have
    $\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]$ $=\prod_{f\in\binom{\{i,j,\ell\}}{2}}(1-p_{f})\prod_{\begin{subarray}{c}(i,u)\\ u\neq j,\ell\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}(j,v)\\ v\neq i,\ell\end{subarray}}(1-p_{jv})\prod_{\begin{subarray}{c}(\ell,w)\\ w\neq i,j\end{subarray}}(1-p_{\ell w})$
    $\leq e^{-d_{i}-d_{j}-d_{\ell}+p_{ij}+p_{i\ell}+p_{j\ell}}\leq e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]$ (36)
  • Case $3$: $i\neq j\neq k\neq\ell$. Notice that
    $\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]=\mathbb{E}[\widetilde{Z}_{i}]\cdot\mathbb{E}[\widetilde{Z}_{j}]\cdot\mathbb{E}[\widetilde{Z}_{k}]\cdot\mathbb{E}[\widetilde{Z}_{\ell}]=e^{-d_{i}-d_{j}-d_{k}-d_{\ell}}.$
    Therefore, similar to case $1$, we have
    $\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]$ $=\prod_{f\in\binom{\{i,j,k,\ell\}}{2}}\!\!\!(1-p_{f})\prod_{\begin{subarray}{c}(i,u)\\ u\neq j,k,\ell\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}(j,v)\\ v\neq i,k,\ell\end{subarray}}(1-p_{jv})\prod_{\begin{subarray}{c}(k,w)\\ w\neq i,j,\ell\end{subarray}}(1-p_{kw})\prod_{\begin{subarray}{c}(\ell,x)\\ x\neq i,j,k\end{subarray}}(1-p_{\ell x})$
    $\leq e^{-d_{i}-d_{j}-d_{k}-d_{\ell}+p_{ij}+p_{ik}+p_{i\ell}+p_{jk}+p_{j\ell}+p_{k\ell}}\leq e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]$ (37)
Consequently,
$\mathbb{E}[\widetilde{N}^{2}]$ $=\sum_{1\leq i<j\leq n}\sum_{1\leq k<\ell\leq n}\Big{[}e^{-d_{i}-d_{j}}\bm{1}_{\{i=k,j=\ell\}}+e^{-d_{i}-d_{j}-d_{\ell}}\bm{1}_{\{j=k\}}+e^{-d_{i}-d_{j}-d_{k}-d_{\ell}}\bm{1}_{\{i\neq j\neq k\neq\ell\}}\Big{]}$
$\leq\sum_{1\leq i<j\leq n}\sum_{1\leq k<\ell\leq n}\Big{[}e^{-d_{i}-d_{j}}+e^{-d_{i}-d_{j}-d_{\ell}}+e^{-d_{i}-d_{j}-d_{k}-d_{\ell}}\Big{]}\leq\mu^{2}+\mu^{3}+\mu^{4},$ (38)
where $\mu\triangleq\sum_{1\leq i\leq n}e^{-d_{i}}$. Combining equations (33) and (34) with (35), (36), (37) and (38) yields that
$\mathbb{E}[N^{2}]\leq e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{N}^{2}]\leq e^{6p_{\max}}\mu^{2}(\mu^{2}+\mu+1).$ (39)
Bounding the first moment. Notice that
$\mathbb{E}[\widetilde{N}]=\sum_{1\leq i<j\leq n}\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}]=\sum_{1\leq i<j\leq n}e^{-d_{i}-d_{j}}=\frac{1}{2}\left(\mu^{2}-\sum_{i=1}^{n}e^{-2d_{i}}\right)\geq\frac{\mu}{2}\left(\mu-\max_{1\leq i\leq n}e^{-d_{i}}\right),$ (40)
where the last step uses that $e^{-2d_{k}}\leq e^{-d_{k}}\cdot\max_{1\leq i\leq n}e^{-d_{i}}$ for each $k$. On the other hand,
$\mathbb{E}[N]$ $=\sum_{1\leq i<j\leq n}\mathbb{E}\left[Z_{i}Z_{j}\right]=\sum_{1\leq i<j\leq n}\Big{[}(1-p_{ij})\prod_{\begin{subarray}{c}(i,u)\\ u\neq j\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}(j,v)\\ v\neq i\end{subarray}}(1-p_{jv})\Big{]}$
$\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}\sum_{1\leq i<j\leq n}\Big{[}e^{-p_{ij}-p_{ij}^{2}}\prod_{\begin{subarray}{c}(i,u)\\ u\neq j\end{subarray}}e^{-p_{iu}-p_{iu}^{2}}\prod_{\begin{subarray}{c}(j,v)\\ v\neq i\end{subarray}}e^{-p_{jv}-p_{jv}^{2}}\Big{]}$
$\stackrel{{\scriptstyle\text{(b)}}}{{\geq}}e^{-(2n-1)p_{\max}^{2}}\cdot\sum_{1\leq i<j\leq n}e^{-d_{i}-d_{j}+p_{ij}}\stackrel{{\scriptstyle\text{(c)}}}{{\geq}}e^{-(2n-1)p_{\max}^{2}}\cdot\mathbb{E}[\widetilde{N}],$ (41)
where (a) uses that $1-x\geq e^{-x-x^{2}}$ for all $x\in[0,1/2]$, (b) uses that $e^{-p_{ij}^{2}}\geq e^{-p_{\max}^{2}}$, and (c) uses that $e^{p_{ij}}\geq 1$ for all $i$ and $j$. Finally, combining equations (39), (40) and (41), we have that
$\frac{\mathbb{E}[N]^{2}}{\mathbb{E}[N^{2}]}\geq e^{-(4n-2)p_{\max}^{2}-6p_{\max}}\frac{\mathbb{E}[\widetilde{N}]^{2}}{\mathbb{E}[\widetilde{N}^{2}]}\geq\exp\left(-(4n-2)p_{\max}^{2}-6p_{\max}\right)\frac{\mu^{2}-2\mu\max_{1\leq i\leq n}e^{-d_{i}}}{4(\mu^{2}+\mu+1)},$
where the last step uses that
$\mathbb{E}[\widetilde{N}]^{2}\geq\frac{\mu^{2}}{4}\left(\mu^{2}-2\mu\cdot\max_{1\leq i\leq n}e^{-d_{i}}+\max_{1\leq i\leq n}e^{-2d_{i}}\right)\geq\frac{\mu^{2}}{4}\left(\mu^{2}-2\mu\max_{1\leq i\leq n}e^{-d_{i}}\right).$
@@ Appendix B On negative results for exact recovery with two graphs
B.1 Proof of Lemma 9
Proof of Lemma 9.    Let $N$ be the number of pairs of isolated nodes in $H$. It follows from Lemma 29 that
$\mathbb{P}\left(N\geq 1\right)$ $\geq\exp\left(-(2n-4)q_{\max}^{2}-6q_{\max}\right)\cdot\frac{\lambda^{2}-2\lambda\max_{1\leq i\leq n}e^{-r_{i}}}{4(\lambda^{2}+\lambda+1)}$
$\geq(1-o(1))\cdot\frac{\lambda^{2}-2\lambda\max_{1\leq i\leq n}e^{-r_{i}}}{4(\lambda^{2}+\lambda+1)},$ (42)
where $\lambda\triangleq\sum_{i=1}^{n}e^{-r_{i}}$, and the last step in (42) uses that $q_{\max}=o(1/\sqrt{n})$. Let $r_{[1]}\leq r_{[2]}\leq\cdots\leq r_{[n]}$ denote a sorted copy of the sequence $r_{1},\cdots,r_{n}$. Consider the following three cases.
  • Case $1$. $r_{[1]}=\omega(1)$. In this setting, it follows that
    $\lambda\max_{1\leq i\leq n}e^{-r_{i}s(1-(1-s)^{m-1})}=\lambda e^{-r_{[1]}s(1-(1-s)^{m-1})}=o(1),$
    since $\lambda=\Omega(1)$. Thus,
    $\mathbb{P}\left(N\geq 1\right)\geq(1-o(1))\cdot\frac{\lambda^{2}-o(1)}{4(\lambda^{2}+\lambda+1)}=\Omega(1).$
  • Case $2$. $r_{[2]}=O(1)$. In this setting, there exists a constant $C>0$ such that $C\leq e^{-r_{[2]}}\leq e^{-r_{[1]}}$. Let $A$ denote the adjacency matrix of $G$ and let $R_{[1]}$ and $R_{[2]}$ denote the degree of two nodes with mean degree $r_{[1]}$ and $r_{[2]}$ respectively. Then,
    $\mathbb{P}\left(R_{[1]}=0,R_{[2]}=0\right)$ $=\mathbb{P}\left(R_{[1]}=0\right)\cdot\mathbb{P}\left(R_{[2]}=0\mid R_{[1]}=0\right)\geq\mathbb{P}\left(R_{[1]}=0\right)\mathbb{P}\left(R_{[2]}=0\right)$
    $=\prod_{i\neq[1]}(1-q_{[1]i})\prod_{j\neq[2]}(1-q_{[2]j})\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}\prod_{i\neq[1]}e^{-q_{[1]i}-q_{[1]i}^{2}}\prod_{j\neq[2]}e^{-q{[2]j}-q_{[2]j}^{2}}$
    $=e^{-r_{[1]}}e^{-r_{[2]}}\cdot\exp\left(-\sum_{i\neq[1]}q_{[1]i}^{2}-\sum_{j\neq[2]}q_{[2]j}^{2}\right)$
    $\stackrel{{\scriptstyle\text{(b)}}}{{\geq}}\exp\left(-2(n-1)q_{\max}^{2}\right)\cdot C^{2}\stackrel{{\scriptstyle\text{(c)}}}{{=}}\Omega(1).$
    Here, (a) uses the inequality $1-x\geq e^{-x-x^{2}}$ whenever $x\in[0,1/2]$, (b) uses that $e^{-q_{ij}^{2}}\geq e^{-q_{\max}^{2}}$ for all $i,j$ and also that $C\leq e^{-r_{[2]}}\leq e^{-r_{[1]}}$. Finally, (c) uses that $\exp\left(-2(n-1)q_{\max}^{2}\right)=\Omega(1)$ whenever $q_{\max}=o(1/\sqrt{n})$.
  • Case $3$. $r_{[2]}=\omega(1)$. Consider the induced subgraph $G^{\prime}$ of $G$ on the vertex set $V=V-\{[1]\}$. The mean degrees of $G^{\prime}$ satisfy $r_{[i]}^{\prime}\leq r_{[i]}$, and so it follows that
    $\lambda^{\prime}\triangleq\sum_{i=2}^{n}e^{-d^{\prime}_{[i]}}\leq\sum_{i=2}^{n}e^{-r_{[i]}}=\Omega(1),$
    and the smallest mean degree in $G^{\prime}$ diverges, i.e. $r_{[2]}=\omega(1)$. Equivalently, the graph $G^{\prime}$ satisfies the conditions for Case $1$ above, and so there is a constant lower bound $C$ on the probability that the number of pairs of isolated nodes $N^{\prime}$ in $G^{\prime}$ is at least one. If node $[1]$ does not connect to an isolated pair of nodes in $G^{\prime}$, then the pair remains isolated in $G$ as well. Therefore,
    $\mathbb{P}\left(N\geq 1\right)\geq(1-q_{\max})^{2}\cdot\mathbb{P}\left(N^{\prime}\geq 1\right)\geq(1-q_{\max})^{2}\cdot C=\Omega(1).$